OpenStudy (anonymous):

Need some help with this one guys.. Find how many quarts of 6% butterfat milk and 3% butterfat milk should be mixed to yield 75 quarts of 5% butterfat milk

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

75 quarts * .05% butterfat milk = (75-x)(.06) + (75-y)(.03) Letting x be 25 and y be 50 we see that x + y = 75 and 50(.06) +25(.03) = 3 + .75 which equals 3.75....... We then check that 75(.05) = 3.75 which it indeed does.